OpenStudy (anonymous):

can you explain the process to finding the derivative of this? y = (x + 1)^2(x^2 + 1)^ -3

OpenStudy (mimi_x3):

Use the product rule.

OpenStudy (mimi_x3):

u = (x+1)^2 u' = 2(x+1) v = (x^2+1)^-3 v' = -3(x^2+1)^-4 Can you do it now?

OpenStudy (anonymous):

No i got that far but it gets complicated for me....

OpenStudy (mimi_x3):

woops, v' = -6x(x^2+1)^-4

OpenStudy (anonymous):

complicated !!!!! let me do it

OpenStudy (anonymous):

Or use logarithmic differentiation: lny = ln( (x + 1)^2(x^2 + 1)^-3) lny = 2ln(x+1) -3ln(x^2 + 1) y'/y = 2/(x+1) - 6x/(x^2 + 1) y' = [2/(x+1) - 6x/(x^2 + 1)] ((x + 1)^2(x^2 + 1)^-3)

OpenStudy (anonymous):

y = (x + 1)^2(x^2 + 1)^ -3 y' = [(x + 1)^2][-3(x^2+1)^-4]+[(x^2 + 1)^ -3][2(x+1)] = -3 (x+1)^2 2x+2 ---------- + ------------- (x^2+1)^4 (x^2 + 1)^ 3 = -3 (x+1)^2+ (2x+2)(x^2+1) ------------------------ (x^2+1)^4

OpenStudy (anonymous):

logs rule!

OpenStudy (anonymous):

simplify it further you'll be getting the answer

OpenStudy (anonymous):

Dhashni im doing a study for calc and I got up to the part where you have the 3rd equal sign but how do i simplify it to the the forth part? because thats the answer that I needed to get it right. thanks for everyones help. =)